Complementary Events
पूरक घटनाएँ
Lesson language · भाषा
Now try 5 questions on this
Quick practice, straight from the lesson's topic. अब आपकी बारी।
Try 5 questions →
In this lesson
Summary
- The idea
- $P(E) + P(\bar{E}) = 1$, so $P(\bar{E}) = 1 - P(E)$ ($\bar{E}$ = not $E$)
- Example 1
- Example one.
- Two dice
- When two dice are rolled, there are thirty-six possible outcomes.
- Example 2
- P(at least one 6) $= 1 - \dfrac{25}{36} = \dfrac{11}{36}$
- Careful
- First, remember that P of E plus P of not E equals one.
- Your turn
- P(at least one tail) $= 1 - \dfrac{1}{4} = \dfrac{3}{4}$ ✓
हिन्दी में सारांश
- मूल विचार
- $P(E) + P(\bar{E}) = 1$, तो $P(\bar{E}) = 1 - P(E)$ ($\bar{E}$ = $E$ नहीं)
- उदाहरण 1
- उदाहरण एक।
- दो पासे
- जब दो पासे फेंके जाते हैं, तो छत्तीस संभावित परिणाम होते हैं।
- उदाहरण 2
- P(कम से कम एक छह) $= 1 - \dfrac{25}{36} = \dfrac{11}{36}$
- सावधान
- पहली, याद रखिए कि पी ऑफ़ ई प्लस पी ऑफ़ नॉट ई बराबर एक।
- अब आपकी बारी
- P(कम से कम एक पट) $= 1 - \dfrac{1}{4} = \dfrac{3}{4}$ ✓
Save your progress and get the full chapter
Free account: practice sets, Quests, exam papers and your own progress.
Sign up free